OpenStudy (anonymous):

HELP ME. Enter the quadratic equation in standard form. x^2-4x-29=-10

OpenStudy (lgbasallote):

standard form -> ax^2 + bx + c = 0

OpenStudy (anonymous):

add 10

OpenStudy (lgbasallote):

just make the right side zero then you're all set

OpenStudy (lgbasallote):

by doing that

OpenStudy (anonymous):

x^2 - 4x -29 + 10 = 0 => x^2 - 4x - 19.

Parth (parthkohli):

\(\Large \color{purple}{\rightarrow x^2 + 4x + 29 = 10 }\) Multiply both sides by -1 and you get the answer.

Parth (parthkohli):

You can also subtract 29 from both sides.

Parth (parthkohli):

I've done my job @pratu043 @satellite73 @lgbasallote

OpenStudy (anonymous):

Or subtract 10 from both sides.

OpenStudy (lgbasallote):

@ParthKohli you can NOT subtract 29 from both sides...like i stated standard form is ax^2 + bx + c = 0 so the right side MUST be zero...just clearing your misconception

Parth (parthkohli):

So subtract 10?

OpenStudy (lgbasallote):

yup only way

OpenStudy (anonymous):

So the actual answert is?...

Parth (parthkohli):

Use my way, just don't do that 29 thingy. Multiply -1 to both sides.